OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{prove that the equation has no integer solution}\hspace{.33em}\\~\\ x^2+y^2=57\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (fibonaccichick666):

well, hmm let's see

OpenStudy (fibonaccichick666):

So do you agree that if we restrict ourselves to integers only the x^3 could ever be negative?

OpenStudy (fibonaccichick666):

btw, I'm still not sure I buy this claim

OpenStudy (kainui):

When x is positive, we know that both x and y are odd at the very least. The hardest thing to show here looks like when y is really large and x is really large and negative.

OpenStudy (fibonaccichick666):

yea, That's what I was thinking about Kainui, do you think if we did some form of mod or common divisor thing we could exclude large values of each?

OpenStudy (kainui):

Beats me honestly, I'm sorta just stumped trying out random stuff.

OpenStudy (fibonaccichick666):

And I think we could quickly eliminate all positive x,y

OpenStudy (kainui):

It might be nice to rewrite x^3 as -x^3 since it is arbitrary.

OpenStudy (fibonaccichick666):

we'd only need the cubes less than 56 and squares less than 56 to eliminate positive combos

OpenStudy (kainui):

Yeah we can easily exhaust our way through the possibilities there, and if we're proving it I mean we can just do that last since it's straight forward, I am imagining that will work out simply enough but idk. \[\large y^2 = 57+x^3\] That's kind of what I'm looking at atm, idk.

OpenStudy (fibonaccichick666):

well, what if we reduced this to mod19. \(\frac{57}{3}=19\)

OpenStudy (fibonaccichick666):

we would be able to look at only x,y=1,2,3...19. Set's of combinations

OpenStudy (fibonaccichick666):

or mod 3 for that matter..

OpenStudy (fibonaccichick666):

so we assume that \(x^3+y^2\equiv 0 mod 3\)

OpenStudy (fibonaccichick666):

or we assume \(x^3+y^2\equiv 0 mod 19\)

OpenStudy (fibonaccichick666):

We just made two equations with two unknowns? @Kainui what do you think?

OpenStudy (kainui):

I am not really all that good at mods honestly so I don't know what to think, it seems like you're doing something though haha.

OpenStudy (kainui):

I would probably do mod 3 since it looks like that means we have less to check?

OpenStudy (fibonaccichick666):

but that would be an issue if they were both 0 mod 3?

OpenStudy (fibonaccichick666):

grr

OpenStudy (kainui):

\[\Large (3n+a)^3 + (3m+b)^2 \equiv 0 \mod 3 \\ \Large a^3 +b^2 \equiv 0 \mod 3\] But now we're looking at a and b only being 0, 1 or 2 so we can rule out some stuff. Maybe we can then do the same thing with mod 19 and look at the intersection of these two sets?

OpenStudy (fibonaccichick666):

maybe like chinese remainder hmm...

OpenStudy (fibonaccichick666):

@mathmath333 what class is this for?

OpenStudy (fibonaccichick666):

I guess a simpler way just to look at it is to graph it as \[y=+-(57-x^3)^{1/2}\]

OpenStudy (fibonaccichick666):

that should only have at max 3 zeroes right?

OpenStudy (kainui):

Ooooh that seems like that method will work actually.

OpenStudy (fibonaccichick666):

but that's not a "proof"

OpenStudy (fibonaccichick666):

that's my only issue

OpenStudy (fibonaccichick666):

but anyways,im annoyed that the guy who asked the question isn't even online so I'm gonna stop

OpenStudy (kainui):

Well I mean how is it not a proof?

OpenStudy (fibonaccichick666):

a graph is not a proof

OpenStudy (kainui):

Unless it crosses at an integer, which are easily checked and we can tell if the function is always increasing with calculus so we know it won't cross again, that kind of thing, seems like it is proof.

OpenStudy (kainui):

I think you are wrong about what is and isn't a proof.

OpenStudy (fibonaccichick666):

^ that is not a proof either

OpenStudy (kainui):

This isn't much different than the IVT.

OpenStudy (fibonaccichick666):

and no, trust me, I have been marked incorrect enough times to know this is not a proof

OpenStudy (fibonaccichick666):

we would have to write this up in a rigorous way using the ivt

OpenStudy (fibonaccichick666):

and some other stuff

OpenStudy (kainui):

Being marked incorrect is not a logical argument, it's a logical fallacy, appealing to authority.

OpenStudy (fibonaccichick666):

no, being told it is not a proof in analysis courses has made me of the notion, that it isn't

OpenStudy (kainui):

Perhaps your teacher expected you to show something was true through a certain method that you were being tested on, which is different. There's no reason to say that this is illegitimate when it shows what is true.

OpenStudy (fibonaccichick666):

it is not rigorous, hence not a proof. They(multiple professors) don't care what method, it must be rigorous, and this is not.

OpenStudy (fibonaccichick666):

My math department is very very analysis heavy

OpenStudy (kainui):

What is not rigorous about it?

OpenStudy (fibonaccichick666):

The fact that we are using a picture one. 2, no formal write up, 3. still not proved

OpenStudy (fibonaccichick666):

I mean if you do a formal ivt write up, maybe

OpenStudy (kainui):

There is no need for a picture, and even if we were, you are basically slapping Euclid and saying he has no rigor. It can be formally written up.

OpenStudy (fibonaccichick666):

but we haven't done that

OpenStudy (fibonaccichick666):

(but we haven't formally written it up, hence not proof)

OpenStudy (kainui):

Hahaha we don't have to, that's @mathmath333 's job, we have found the way to the answer though, that's all I am worried about.

OpenStudy (fibonaccichick666):

lol

OpenStudy (kainui):

Don't get me wrong though, a geometric proof is still a proof.

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\normalsize \text{lol counter example }\hspace{.33em}\\~\\ 2^3+7^2=57\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (mathmath333):

sry guys the question is this one \(x^2+y^2=57\)